Please help me with this linear equation:Write the equation of a line that is perpendicular to y=-0.3x+6.
Answer by Alan3354(69443)

Please help me with this linear equation:Write the equation of a line that is perpendicular to y=-0.3x+6.
-------------------
The slope of y=-0.3x+6 is -0.3
===============
The slope of lines (an infinite # of them) perpendicular is -1/(-0.3) = 10/3
----
The lines would be:
y = (10/3)x + b where b in the y-intercept and can be any number.